Convert 15.33 grams of KCl to moles of KCI. Molar mass of KCl is 74.55 g/mol


Convert 29.14 grams of KCl to moles of KCI. Molar mass of KCl is 74.55g / mol 


Convert 44.33 grams of KCl to moles of LiF. Molar mass of LiF is 25.94 g / mol 


Convert 0.6047 moles of PCl5 to g of PCl5. Molar mass of PCl5 is 208.24 g / mol


Convert 0.8173 moles of BH3 to g of BH3. Molar mass of BH3 is 13.84 g / mol


Convert 0.8912 moles of BH3 to g of BH3. Molar mass of BH3 is 13.84 g / mol




1
Expert's answer
2021-02-26T06:00:18-0500

a) "Molar Mass of KCL=74.55\frac{g}{mol}"


"=\frac{15.34g×mol of KCL} {74.55 g/mol}"


=0.2056moles of KCl


b) "Molar Mass of KCL=74.55\frac{g}{mol}"


"=\frac{29.14g×mol of KCL} {74.55g/mol}"


=0.3909 moles of KCl


c) "Molar Mass of LiF=25.94\frac{g}{mol}"


"=\frac{44.33g×mol of LiF } {25.94/mol}"


=1.7089moles of LiF


d) "Molar Mass of PCl5 =208.24\frac{g}{mol}"


"Grams of PCl5=0.6047 moles of PCl5 \u00d7208.24g\/mol"


=125.92g of "PCl_5"


e) "Molar Mass of BH3=13.84\frac{g}{mol}"


"=Grams of BH3=0.8173moles of BH3 \u00d713.84g\/mol"


=11.31g of "BH_3"


f) "Molar Mass of BH3=13.84 \frac{g}{mol}"

"=Grams of BH3=0.8912moles of BH3 \u00d713.84g\/mol"


=12.33 g of "BH_3"